Jodie bicycles 8 km/h faster than Carlos. In the same time it takes Carlos to bicycle 60 km, Jodie can bicycle 84 km. How fast does each bicyclist travel?
Carlos travels ____ km/h
Jodie travels ____ km/h

Answer by mananth(16946)

Jodie bicycles 8 km/h faster than Carlos. In the same time it takes Carlos to bicycle 60 km, Jodie can bicycle 84 km. How fast does each bicyclist travel?
let Carlos speed at x kmph
Jodie speed = x+8 kmph
..
Carlos distance = 60 km
Jodie distance = 84 km
..
Carlos time = 60/x
Jodie time = 84/(x+8)
..
60/x = 84/(x+8)
84x=60(x+8)
84x=60x+480
84x-60x=480
24x=480
x= 20 kmph
x+8 = 20 +8 = 28 kmph
Carlos travels _20___ km/h
Jodie travels 28____ km/h